Four times the sum of the first and third of three consecutive integers is one less than seven times the middle integer. Find the integer.
Answer by phillv(13)

Let the first integer =X
Therefor the second integer =X+1
and the third integer = X+2
The resulting equation is: 4(X+(X+1))=7(X+2)-1
4(2X+1)=7X+14-1
8X+4=7X+13
8X-7X=13-4
X=9
The first integer is X=9
The second integer is X+1=10
The third integer is X+2=11